Collins Inc has 5,000,000 shares of common stock outstanding with a market price of $9.00 per share. It has 25,000 bonds outstanding, each selling for $1,100. The bonds mature in 12 years, have a coupon rate of 8.5%, and pay coupons annually. The firm's beta is 1.2, the risk free rate is 5%, and the market risk premium is 9%. The tax rate is 35%. Calculate the WACC.
Answer #1

Market value of common stock = 5,000,000 * 9 = 45,000,000

Market value of debt = 25,000 * 1,100 = 27,500,000

Total market value of capital structure = 45,000,000 + 27,500,000 = 72,500,000

Yield to maturity = 7.2258%

Assuming face value to be $1000

Coupon = 8.5% of 1000 = 85

Keys to use in a financial calculator:

FV 1000

PMT 85

N 12

PV -1100

CPT I/Y

Cost of equity = Risk free rate + beta(market risk premium)

Cost of equity = 5% + 1.2(9%)

Cost of equity = 15.80%

WACC = Weights * costs

WACC = (45,000,000 / 72,500,000)*0.158 + (27,500,000 / 72,500,000)*0.072258*(1 - 0.35)

WACC = 0.098069 + 0.017815

WACC = 0.1159 or 11.59%